OpenStudy (anonymous):

In the below system, solve for y in the first equation. 2x + y = 11 5x – 2y = –1

OpenStudy (anonymous):

2x+y=11 5x-2y=-1 -> x=(-2y-1)/5 use this and plug in for x in the first equation 2((-2y-1)/5)+y=11  -4y-2+5y=55 -2+y=55 y=57